Wavelength of red light is 700.5nm what’s the frequency?
1
Expert's answer
2020-03-27T09:01:03-0400

428 ТHz

ν = с/λ, where

λ = wavelength, m

ν = frequency, Hz

c = speed of light

c = 3*108 m/s

λ = 700.5 nm = 700.5*10-9 m

then ν = 3*108/700.5*10-9 = 4.28*1014 Hz = 428*1012 Hz = 428 THz